find two numbers whose sum is 54 and whose difference is 28.
Answer by checkley79(3341)

X+Y=54
X-Y=28 ADD THESE 2 EQUATIONS.
--------------
2X=82
X=82/2
X=41 ANS. FOR THE LARGER NUMBER.
41+Y=54
Y=54-41
Y=13 ANS FOR THE SMALLER NUMBER.
PROOF:
41-13=28
28=28